How many gallons of a 14%-salt solution must be mixed with 6 gallons of a 21%-salt solution to obtain a 17%-salt solution?
Found 2 solutions by robertb, Tatiana_Stebko:
Answer by robertb(5830)

0.14x + 0.21*6 = 0.17(x + 6)
==> 0.14x + 1.26 = 0.17x +1.02
==> -0.03x = -0.24
==> x = 8 gallons of the 14% salt solution

Answer by Tatiana_Stebko(1539)

If x gallons of a 14%-salt
then obtain (x+6) gallons of a 17%-salt
x gallons of a 14%-salt contains 0.14*x gallons of solt
6 gallons of a 21%-salt contains 0.21*6 gallons of solt
(x+6) gallons of a 17%-salt contains 0.17*(x+6) gallons of solt
So equation:
0.14x+0.21*6=0.17*(x+6)
0.14x+1.26=0.17x+1.02
0.03x=0.24
x=8 gallons of a 14%-salt